Find all solutions in [0,2π) to the equation

2sin2(A)=cos(A)+1

SHOW ALL WORK


1
Expert's answer
2020-07-13T19:14:01-0400

Solution.

2sin2(A) = cos(A) + 1.

We know, that sin2x + cos2x = 1 (1)

By subtracting cos2x from both sides of equation (1) we will receive:

sin2x = 1 - cos2x (2)

Now we use equation (2) for our condition of the task, we will receive:

2(1 - cos2(A)) = cos(A) + 1,

Now we open parentheses and subtract cos(A) + 1 from both sides of the equation:

2 - 2cos2(A) - cos(A) - 1 = 0,

-2cos2(A) - cos(A) + 1 = 0

Multiplying both sides by (-1) we will get:

2cos2(A) + cos(A) - 1 = 0.

Let cos(A) = x, then we will get:

2x2 + x - 1 = 0.

Here we have quadratic equation. To find x we will use next formulas:

if ax2 + bx + c = 0, then

D = b2 - 4ac,

"x_1=\frac{\left(- b - \sqrt{D}\right)}{2a}"


"x_2=\frac{\left(-b+\sqrt{D}\right)}{2a}"


Let's find x from our equation:

2x2 + x - 1 = 0

D = 12 - 4 * 2 *(-1) = 1 + 8 = 9 = 32,

x1 = (-1 - 3) / 2 * 2 = -4 / 4 = -1,

x2 = (-1 + 3) / 2 * 2 = 2 / 4 = 1/2.

Now we use our substitution cos(A) = x and we will receive:

cos(A1) = -1 and cos(A2) = 1/2.

A1 = cos-1(-1) = π,

A2 = cos-1(1/2), in [0,2π) we have two values of A: A2 = π/3, A3 = 5π/3.


Answer: A1 = π, A2 = π/3, A3 = 5π/3.